# Studying magnetic circular vortex dichroism effect for photoionization of Rydberg atoms with vortex photons

### v1

@@ Abstract
Rydberg atoms, renowned for their exceptional quantum properties, hold significant importance in quantum physics. The photoionization of Rydberg atoms serves as a critical tool for probing their unique characteristics. In this work, we investigate the photoionization dynamics of Rydberg atoms interacting with vortex photons—a class of structured light carrying intrinsic orbital angular momentum. This interaction gives rise to novel quantum phenomena distinct from conventional photoionization processes. Our results reveal that vortex photons exhibit exceptional sensitivity to the magnetic moments of Rydberg atoms, positioning them as a powerful spectroscopic tool for investigating Rydberg magnetism. It is also demonstrated that the initial photon energy must be carefully selected to observe significant experimental results. Furthermore, the photoionization process displays strong angular momentum selectivity, preferentially favoring configurations where the photon total angular momentum and atomic magnetic moment are aligned. This pronounced asymmetry directly manifests the chiral nature of the vortex photon-Rydberg atom interaction.
@@ I Introduction
Magnetic dichroism effects, which put materials in magnetic fields and is the magnetic-field-induced counterparts of conventional dichroism effects, provide powerful tools for investigating magnetic properties. While magnetic circular dichroism (MCD)—characterizing the different absorption of left- and right-handed circularly polarized light in magnetic fields—is well established [24, 25, 26], magnetic vortex dichroism (MVD) which probes the dichroic response to vortex beams carrying opposite orbital angular momentum exhibits significantly stronger signals than MCD [27, 28, 29].
In this work, we investigate the magnetic circular vortex dichroism (MCVD) effect by comparing the response of polarized Rydberg atoms to vortex photon beams with opposite topological charges and helicities. Through parity analysis, this comparison is shown to be equivalent to reversing the atoms’ magnetic quantum numbers while keeping the photon beams unchanged. We theoretically calculate the photoionization cross-sections of Rydberg atoms interacting with vortex photons and present MCVD results for various initial topological charges and helicities. Specifically, initial vortex photon beam with energy $100$ eV is selected according to experimental condition [30, 31, 32]. The observed dependence of MCVD on the magnetic quantum number demonstrates the sensitivity of vortex photons to the magnetic moment of Rydberg atoms. Furthermore, asymptotic behavior of MCVD for vortex photon with large total angular momentum (TAM) shows angular momentum selectivity of the interaction.
@@ II Theoretical calculations
II.1 Transition matrix for Rydberg photoionization with initial plane wave photon
Under the first Born approximation, the single-electron photoionization matrix element for an atomic state with quantum numbers $(n,l,m_{l})$ by a plane wave photon in coordinate space is given by:
@@ II.1 Transition matrix for Rydberg photoionization with initial plane wave photon
where:
  • $e$ is the electron charge
  • $m_{e}$ is the electron mass
  • $\mathbf{r}$ is the position coordinate
  • $\mathbf{k}$ is the incident photon momentum
  • $\mathbf{\epsilon}_{\mathbf{k}{\Lambda_{\gamma}}}$ is the photon polarization vector (helicity $\Lambda_{\gamma}$})
  • $\mathbf{p}$ is the outgoing electron momentum
  • $\hat{\mathbf{p}}=-i\mathbf{\nabla}$ is the momentum operator
  • $\psi_{f}$ is the final electron state
  • $\psi_{i}^{nlm_{l}}$ is the initial bound state
In spherical coordinates, the momenta are parameterized as $\mathbf{k}=(k,\theta_{\mathbf{k}},\phi_{\mathbf{k}})$ and $\mathbf{p}=(p,\theta_{\mathbf{p}},\phi_{\mathbf{p}})$. The photon polarization vector in Coulomb gauge takes the explicit form:
$\mathbf{\epsilon}_{\mathbf{k}{\Lambda_{\gamma}}}=\begin{pmatrix}-\dfrac{{\Lambda_{\gamma}}}{\sqrt{2}}e^{-i{\Lambda_{\gamma}}\phi_{\mathbf{k}}}\cos^{2}\dfrac{\theta_{\mathbf{k}}}{2}+\dfrac{{\Lambda_{\gamma}}}{\sqrt{2}}e^{i{\Lambda_{\gamma}}\phi_{\mathbf{k}}}\sin^{2}\dfrac{\theta_{\mathbf{k}}}{2}\\[10.0pt] -\dfrac{i}{\sqrt{2}}e^{-i{\Lambda_{\gamma}}\phi_{\mathbf{k}}}\cos^{2}\dfrac{\theta_{\mathbf{k}}}{2}-\dfrac{i}{\sqrt{2}}e^{i{\Lambda_{\gamma}}\phi_{\mathbf{k}}}\sin^{2}\dfrac{\theta_{\mathbf{k}}}{2}\\[10.0pt] \dfrac{{\Lambda_{\gamma}}}{\sqrt{2}}\sin\theta_{\mathbf{k}}\end{pmatrix}$. (2)
@@
${\cal M}_{fi}^{P}=-\frac{e}{m_{e}}(\mathbf{\epsilon}_{\mathbf{k}{\Lambda_{\gamma}}}\cdot\mathbf{q})\tilde{\psi}_{i}^{nlm}(\mathbf{q}),$ (3)
@@
The atomic wavefunction in momentum space has the closed-form expression [33]:
$\tilde{\psi}_{nlm_{l}}(\mathbf{q})$ $=2^{2l+2}l!(-i)^{l}n^{-l-2}\sqrt{\frac{2(n-l-1)!}{\pi Z^{3}(n+l)!}}\frac{Z^{2l+4}}{(q^{2}a^{2}+\delta^{2})^{l+2}}$
$\quad\times C_{n-l-1}^{l+1}\left(\frac{q^{2}a^{2}-\delta^{2}}{q^{2}a^{2}+\delta^{2}}\right)\left(\frac{qa}{Z}\right)^{l}Y_{l}^{m_{l}}(\theta_{\mathbf{q}},\phi_{\mathbf{q}}),$ (5)
with:
  • $Z$: nuclear charge
  • $a$: Bohr radius
  • $\delta\equiv Z/n$
  • $C_{n}^{\alpha}(x)$: Gegenbauer polynomial
  • $Y_{l}^{m_{l}}$: spherical harmonic function
II.2 Transition matrix for Rydberg photoionization with initial vortex photon
To demonstrate our key physical results, we focus specifically on alkali Rydberg atoms with the principal quantum number $n=l+1$. Single electron ionization is considered and thus $Z$ is set to be one.
II.2.1 Momentum Space Wavefunction for the selected Rydberg atoms
The atomic wavefunction in momentum space is given by:
$\tilde{\psi}_{nl{m_{l}}}(\mathbf{q})$ $=2^{2n}l!(-i)^{l}n^{-l-2}\sqrt{\frac{2}{\pi(2l+1)!}}\frac{q^{l}a^{l}Y_{l}^{m_{l}}(\theta_{\mathbf{q}},\phi_{\mathbf{q}})}{(q^{2}a^{2}+\delta^{2})^{l+2}}$
@@ II.2.1 Momentum Space Wavefunction for the selected Rydberg atoms
$\hat{N}=2^{2n}l!(-i)^{l}n^{l+2}a^{l}\sqrt{\frac{2}{\pi(2l+1)!}},$ (7)
and $\delta=1/n$ for Rydberg alkali atoms.
II.2.2 Plane Wave Photoionization for the selected Rydberg atoms
The transition amplitude for plane wave photons is:
@@ II.2.2 Plane Wave Photoionization for the selected Rydberg atoms
II.2.3 Vortex Photon Photoionization for the selected Rydberg atoms
The most convenient vortex states are the Bessel vortex states which will be used in this paper. A Bessel vortex state for photon is
where $N_{B}=(-i)^{m_{\gamma}}\sqrt{{\kappa}/{2\pi}}$, $\vec{k}_{\perp}$ is transverse part of momentum $\mathbf{k}$ (longitudinal part is written as $k_{z}$). In spherical coordinates, the momenta are parameterized as $\mathbf{k}=(k,\theta_{\mathbf{k}},\phi_{\mathbf{k}})$. There are four quantum numbers: energy $\omega$, transverse momentum $\kappa$, total angular momentum (TAM) quantum number $m_{\gamma}$ and helicity $\Lambda_{\gamma}$}. The state has TAM $\hbar m_{\gamma}$.
The position-dependent photon current density is [34] :
where $r_{\perp}$ denotes the transverse coordinate and $J_{n}(x)$ is the first kind Bessel function.
@@ II.2.3 Vortex Photon Photoionization for the selected Rydberg atoms
${\cal M}^{V}_{fi}$ $=N_{B}\int_{0}^{2\pi}\frac{d\phi_{\mathbf{k}}}{2\pi}{\rm e}^{im_{\gamma}\phi_{\mathbf{k}}}M^{P}_{fi}(\mathbf{k},\mathbf{p})$
@@
$R_{l}^{m_{l}}(\mathbf{r}+\mathbf{a})$ $=\sum_{\sigma=0}^{l}\sum_{\mu=-\sigma}^{\sigma}{\sqrt{\binom{2l}{2\sigma}}}R_{\sigma}^{\mu}(\mathbf{r})R_{l-\sigma}^{{m_{l}}-\mu}(\mathbf{a})$
@@
where $\binom{2l}{2\sigma}$} is the binomial coefficient, we obtain the expanded formula:
@@
$\quad\times\sum_{\sigma=0}^{l}\sum_{\mu=\mu_{\rm min}}^{\mu_{\rm max}}f_{l{m_{l}},\sigma\mu}(\mathbf{p})e^{i({m_{l}}-\mu)\phi},$ (15)
@@
$f_{l{m_{l}},\sigma\mu}(\mathbf{p})$ $=\sqrt{\frac{(2l+1)(\sigma-\mu)!(l-\sigma-{m_{l}}+\mu)!}{4\pi(\sigma+\mu)!(l-\sigma+{m_{l}}-\mu)!}\binom{2l}{2\sigma}}$
$\quad\times p^{\sigma}(-k)^{l-\sigma}P_{\sigma}^{\mu}(\cos\theta_{\mathbf{p}})P_{l-\sigma}^{{m_{l}}-\mu}(\cos\theta_{\mathbf{k}})$
@@
$P_{\sigma}^{\mu}$ is the associated Laguerre polynomial and the limitation of $\mu$ is: $\mu_{\rm min}={\rm Max}(-l+{m_{l}}+\sigma,\,-\sigma)$ and $\mu_{\rm max}={\rm Min}(l+{m_{l}}-\sigma,\,\sigma)$.
After substituting Eq.II.2.3 into Eq.II.2.3 and performing the variable change $z=e^{i\phi_{\mathbf{k}}}$, we obtain:
${\cal M}_{fi}^{V}$ $=\frac{i(-1)^{n}N_{B}N_{P}}{2\pi(na)^{2l+4}}e^{i({m_{l}}+m_{\gamma})\phi_{\mathbf{p}}}\sum_{\sigma=0}^{l}\sum_{\mu=\mu_{\rm min}}^{\mu_{\rm max}}f_{l{m_{l}},\sigma\mu}(\mathbf{p})$
@@
The contour integral in Eq. (II.2.3) can be evaluated using the residue theorem. The explicit analytical result is presented in Appendix A.
II.3 Differential Cross Section

### v2

@@ Abstract
Rydberg atoms, renowned for their exceptional quantum properties, hold significant importance in quantum physics. The photoionization of Rydberg atoms serves as a critical tool for probing their unique characteristics. In this work, we investigate the photoionization dynamics of hydrogen-like Rydberg alkali atoms interacting with vortex photons—a class of structured light carrying intrinsic orbital angular momentum. This process gives rise to novel quantum phenomena distinct from conventional photoionization processes. Our results reveal that vortex photons exhibit exceptional sensitivity to the magnetic moments of Rydberg atoms, positioning them as a powerful spectroscopic tool for investigating Rydberg magnetism. It is also demonstrated that the initial photon energy must be carefully selected to observe significant experimental results. Furthermore, the photoionization process displays strong angular momentum selectivity, preferentially favoring configurations where the photon total angular momentum and atomic magnetic moment are aligned. This pronounced asymmetry directly manifests the chiral nature of the vortex photon-Rydberg atom collisions.
@@ I Introduction
Magnetic dichroism effects, which put materials in magnetic fields and is the magnetic-field-induced counterparts of conventional dichroism effects, provide powerful tools for investigating magnetic properties. Magnetic circular dichroism (MCD) effects, which characterize the different absorption of left- and right-handed circularly polarized light for material in magnetic fields, is well established as powerful tool for magnetic property detection [25, 31, 32]. Magnetic vortex dichroism (MVD) effects, which probe the dichroic response of matter to vortex beams carrying opposite OAM, have been proven to exhibit significantly stronger signals than MCD [33, 1, 11]. Based on the extension of the conventional dichroism effect to CVD when both spin and orbital angular momentum are considered, we propose the magnetic circular vortex dichroism (MCVD) effect for magnetic systems. This phenomenon has not been investigated to date.
The collision between polarized vortex photons and polarized Rydberg atoms provides an ideal platform for studying the MCVD effect. On the one hand, polarized vortex photons have been successfully generated and utilized in atomic photo-excitation processes [30]. Polarized vortex states can be precisely controlled, and the atomic targets can be localized with sub-nanometer precision using optical trapping techniques. On the other hand, the production and localization of polarized Rydberg atoms have also been achieved [27]. These atoms can be confined in carefully designed optical tweezers, with their polarization direction defined by a weak static magnetic field. Given these advancements, it is now feasible to design experiments to study semi-central collisions between vortex photons and single Rydberg atom target.
In this work, we investigate the MCVD effect by comparing the response of polarized Rydberg atoms to vortex photon beams with opposite topological charges and helicities. Through parity analysis, this comparison is shown to be equivalent to reversing the atoms’ magnetic quantum numbers while keeping the photon beams unchanged. We theoretically calculate the photoionization cross-sections of Rydberg atoms interacting with vortex photons and present MCVD results for various initial topological charges and helicities. Specifically, initial vortex photon beam with energy $100$ eV is selected according to experimental condition [7, 18, 22]. The observed dependence of MCVD on the magnetic quantum number demonstrates the sensitivity of vortex photons to the magnetic moment of Rydberg atoms. Furthermore, asymptotic behavior of MCVD for vortex photon with large total angular momentum (TAM) shows angular momentum selectivity of the process.
@@ II Theoretical calculations
In this paper, the target is one single hydrogen-like Rydberg alkali atom. We investigate photoionization of its valence electron and use hydrogen-like wave function for calculations. As being discussed in the introduction section, semi-central collision with such single atom target is feasible nowadays. We begin with calculation of initial plane wave photon case and then go to initial vortex photon case.
II.1 Vortex states and collision scenario
Prior to undertaking the calculations, it is instructive to briefly introduce the vortex states and the collision scenario that will be examined.
II.1.1 Vortex states
A vortex state typically refers to a type of structured wave that carries intrinsic OAM. It features a singularity line where the amplitude vanishes and the phase becomes indeterminate. The probability current of such a state flows helically around this singularity line, providing a visual representation of its intrinsic OAM.
For our purposes, the most convenient type is the Bessel vortex state, which will be used throughout this paper. A Bessel vortex state for a photon is given by [4]
where $N_{B}=(-i)^{m_{\gamma}}\sqrt{{\kappa}/{2\pi}}$ is the normalization factor, and $\vec{k}_{\perp}$ represents the transverse part of the momentum $\mathbf{k}$ (its longitudinal part is denoted $k_{z}$). In spherical coordinates, the momentum is parameterized as $\mathbf{k}=(k,\theta_{\mathbf{k}},\phi_{\mathbf{k}})$. The state is characterized by four quantum numbers: the energy $\omega$, the transverse momentum $\kappa$, the total angular momentum (TAM) projection $m_{\gamma}$, and the helicity $\Lambda_{\gamma}$}. It carries a TAM of $\hbar m_{\gamma}$, where $m_{\gamma}$ can be any integer. The quantity $\bar{m}_{\gamma}=m_{\gamma}-\Lambda_{\gamma}$ is defined as the topological charge, which dictates the phase variation of the state. The transverse intensity profile consists of bright and dark concentric rings with a central null point. An example for a Bessel vortex state is shown in Fig. 1.
Figure 1: Transverse intensity profile of a Bessel vortex photon with $\omega=100$ eV, $\theta_{\mathbf{k}}=\pi/10$, $m_{\gamma}=1$, and $\Lambda_{\gamma}=-1$. All transverse distances are in nanometers.
This state is a superposition of multiple plane wave components. In momentum space, these components are distributed on a circle orthogonal to the propagation direction, and their momentum vectors collectively form a circular cone, as depicted in Fig. 2.
Figure 2: Momentum distribution of a monochromatic Bessel vortex state with topological charge $\bar{m}_{\gamma}=1$. The black circle of radius $\kappa$ indicates the distribution in momentum space. All momentum vectors lie on a circular cone. The phase factor of the plane wave components advances by $2\pi\bar{m}_{\gamma}$ around the circle (phases are color-coded). Here, $k_{z}$ is the longitudinal momentum and $\kappa$ is the fixed transverse momentum magnitude for all components.
Finally, the position-dependent photon current density is given by [6]:
where $r_{\perp}$ denotes the transverse coordinate and $J_{n}(x)$ is the Bessel function of the first kind.
II.1.2 Collision scenario
For a central collision between a vortex photon and a Rydberg atom, the scattering geometry is illustrated in Fig. 3.
Figure 3: Geometry of a central collision between a vortex photon and a Rydberg atom. The polarized Rydberg atom is located within a weak, constant magnetic field. The incident vortex photon (represented by the blue cone with cone angle $\theta_{k}$) propagates along an axis collinear with the magnetic field direction. Here, $k_{z}$ and $\kappa$ denote the photon’s longitudinal momentum and transverse momentum modulus, respectively. The atom is positioned directly on the singularity line of the vortex photon.
In this setup, the target atom is placed within a weak, constant magnetic field that defines its quantization axis. A vortex photon is incident on the atom, with the atom located precisely on the singularity line of the vortex beam. The direction of the magnetic field and the propagation direction of the vortex photon are collinear.
II.2 Transition matrix for Rydberg photoionization with initial plane wave photon
Under the first Born approximation, the single-electron photoionization matrix element for an atomic state by a plane wave photon in coordinate space is given by:
@@ II.2 Transition matrix for Rydberg photoionization with initial plane wave photon
where $e$ is the electron charge, $m_{e}$ is the electron mass, $\mathbf{r}$ is the position coordinate, $\mathbf{k}$ is the incident photon momentum, $\mathbf{\epsilon}_{\mathbf{k}{\Lambda_{\gamma}}}$ is the photon polarization vector with helicity $\Lambda_{\gamma}$}, $\mathbf{p}$ is the outgoing electron momentum, $\hat{\mathbf{p}}=-i\mathbf{\nabla}$ is the momentum operator, $\psi_{f}$ is the final electron state and $\psi_{nlm_{l}}$ is the initial bound state. $(n,l,m_{l})$ are principle quantum number, orbital quantum number and magnetic quantum number of atomic state respectively.
In spherical coordinates, the momenta are parameterized as $\mathbf{k}=(k,\theta_{\mathbf{k}},\phi_{\mathbf{k}})$ and $\mathbf{p}=(p,\theta_{\mathbf{p}},\phi_{\mathbf{p}})$. The photon polarization vector in Coulomb gauge takes the explicit form [4]:
$\mathbf{\epsilon}_{\mathbf{k}{\Lambda_{\gamma}}}=\begin{pmatrix}-\dfrac{{\Lambda_{\gamma}}}{\sqrt{2}}e^{-i{\Lambda_{\gamma}}\phi_{\mathbf{k}}}\cos^{2}\dfrac{\theta_{\mathbf{k}}}{2}+\dfrac{{\Lambda_{\gamma}}}{\sqrt{2}}e^{i{\Lambda_{\gamma}}\phi_{\mathbf{k}}}\sin^{2}\dfrac{\theta_{\mathbf{k}}}{2}\\[10.0pt] -\dfrac{i}{\sqrt{2}}e^{-i{\Lambda_{\gamma}}\phi_{\mathbf{k}}}\cos^{2}\dfrac{\theta_{\mathbf{k}}}{2}-\dfrac{i}{\sqrt{2}}e^{i{\Lambda_{\gamma}}\phi_{\mathbf{k}}}\sin^{2}\dfrac{\theta_{\mathbf{k}}}{2}\\[10.0pt] \dfrac{{\Lambda_{\gamma}}}{\sqrt{2}}\sin\theta_{\mathbf{k}}\end{pmatrix}$, (4)
where $\Lambda_{\gamma}$ is helicity of the state.
@@
${\cal M}_{fi}^{P}(\mathbf{k},\mathbf{p})=-\frac{e}{m_{e}}(\mathbf{\epsilon}_{\mathbf{k}{\Lambda_{\gamma}}}\cdot\mathbf{q})\tilde{\psi}_{i}^{nlm}(\mathbf{q}),$} (5)
@@
We investigate the photoionization of the valence electron in alkali metal atoms, where only the outermost electron is ionized. Due to their hydrogen-like character, the ionization process can be described using hydrogen atom wave functions. The hydrogen-like atomic wavefunction in coordinate space is
${\psi}_{nlm_{l}}(\mathbf{r})$ $=\sqrt{\Big(\frac{2}{na}\Big)^{3}\frac{(n-l-1)!}{2n(n+l)!}}{\rm e}^{-r/na}\Big(\frac{2r}{na}\Big)^{l}$
$\quad\times\Big[L_{n-l-1}^{2l+1}\Big(\frac{2r}{na}\Big)\Big]Y_{l}^{m_{l}}(\theta,\phi),$ (7)
where $a$ is the Bohr radius, $(r,\theta,\phi)$ are spherical coordinates, $L_{\alpha}^{\beta}$ is the associated Laguerre polynomials and $Y_{l}^{m_{l}}$ is the spherical harmonic function. The atom’s polarization axis has been chosen to be $z$ axis. By Fourier transform, we get its formula in momentum space[26]:
$\tilde{\psi}_{nlm_{l}}(\mathbf{q})$ $=2^{2l+2}l!(-i)^{l}n^{-l-2}\sqrt{\frac{2(n-l-1)!}{\pi(n+l)!}}\frac{1}{(q^{2}a^{2}+\delta^{2})^{l+2}}$
$\quad\times C_{n-l-1}^{l+1}\left(\frac{q^{2}a^{2}-\delta^{2}}{q^{2}a^{2}+\delta^{2}}\right)({qa})^{l}Y_{l}^{m_{l}}(\theta_{\mathbf{q}},\phi_{\mathbf{q}}),$ (8)
where $\delta=1/n$ and $C_{n}^{\alpha}(x)$ is the Gegenbauer polynomial.
II.3 Transition matrix for Rydberg photoionization with initial vortex photon
II.3.1 Momentum Space Wavefunction for the selected Rydberg atoms
For simplicity, we choose $n=l+1$. The atomic wavefunction in momentum space is given by:
$\tilde{\psi}_{n{{(n-1)}}{m_{l}}}(\mathbf{q})$ $=2^{2n}l!(-i)^{l}n^{-l-2}\sqrt{\frac{2}{\pi(2l+1)!}}$
$\quad\times\frac{q^{l}a^{l}Y_{l}^{m_{l}}(\theta_{\mathbf{q}},\phi_{\mathbf{q}})}{(q^{2}a^{2}+\delta^{2})^{l+2}}$
@@ II.3.1 Momentum Space Wavefunction for the selected Rydberg atoms
$\hat{N}=2^{2n}l!(-i)^{l}n^{l+2}a^{l}\sqrt{\frac{2}{\pi(2l+1)!}}.$ (10)
II.3.2 Plane Wave Photoionization for the selected Rydberg atoms
According to Eq. 5, Eq. II.2 and Eq. II.3.1, the transition amplitude for plane wave photons is:
@@ II.3.2 Plane Wave Photoionization for the selected Rydberg atoms
II.3.3 Vortex Photon Photoionization for the selected Rydberg atoms
@@
${\cal M}^{V}_{fi}(\mathbf{p})$ $=N_{B}\int_{0}^{2\pi}\frac{d\phi_{\mathbf{k}}}{2\pi}{\rm e}^{im_{\gamma}\phi_{\mathbf{k}}}M^{P}_{fi}(\mathbf{k},\mathbf{p})$
@@
$R_{l}^{m_{l}}(\mathbf{r}+\mathbf{a})$ $=\sum_{{{\rho}}=0}^{l}\sum_{\mu=-{{\rho}}}^{{{\rho}}}{\sqrt{\binom{2l}{2{{\rho}}}}}R_{{{\rho}}}^{\mu}(\mathbf{r})R_{l-{{\rho}}}^{{m_{l}}-\mu}(\mathbf{a})$
@@
where $\binom{2l}{2{{\rho}}}$} is the binomial coefficient, we obtain the expanded formula:
@@
$\quad\times\sum_{{{\rho}}=0}^{l}\sum_{\mu=\mu_{\rm min}}^{\mu_{\rm max}}f_{l{m_{l}},{{\rho}}\mu}(\mathbf{p})e^{i({m_{l}}-\mu)\phi},$ (16)
@@
$f_{l{m_{l}},{{\rho}}\mu}(\mathbf{p})$ $=\sqrt{\frac{(2l+1)({{\rho}}-\mu)!(l-{{\rho}}-{m_{l}}+\mu)!}{4\pi({{\rho}}+\mu)!(l-{{\rho}}+{m_{l}}-\mu)!}\binom{2l}{2{{\rho}}}}$
$\quad\times p^{{{\rho}}}(-k)^{l-{{\rho}}}P_{{{\rho}}}^{\mu}(\cos\theta_{\mathbf{p}})P_{l-{{\rho}}}^{{m_{l}}-\mu}(\cos\theta_{\mathbf{k}})$
@@
$P_{{{\rho}}}^{\mu}$ is the associated Laguerre polynomial and the limitation of $\mu$ is: $\mu_{\rm min}={\rm Max}(-l+{m_{l}}+{{\rho}},\,-{{\rho}})$ and $\mu_{\rm max}={\rm Min}(l+{m_{l}}-{{\rho}},\,{{\rho}})$.
After substituting Eq.II.3.3 into Eq.II.3.3 and performing the variable change $z=e^{i\phi_{\mathbf{k}}}$, we obtain:
${\cal M}_{fi}^{V}(\mathbf{p})$ $=\frac{i(-1)^{n}N_{B}N_{P}}{2\pi(na)^{2l+4}}e^{i({m_{l}}+m_{\gamma})\phi_{\mathbf{p}}}\sum_{{{\rho}}=0}^{l}\sum_{\mu=\mu_{\rm min}}^{\mu_{\rm max}}f_{l{m_{l}},{{\rho}}\mu}(\mathbf{p})$
@@
The contour integral in Eq. (II.3.3) can be evaluated using the residue theorem. The explicit analytical result is presented in Appendix A.
II.4 Differential Cross Section